# DBLP: Noise Bridge Consistency Distillation For Efficient And Reliable Adversarial Purification

### v1

@@ Introduction
Deep neural networks (DNNs) have achieved remarkable success across a wide range of tasks in recent years. However, their widespread deployment has raised increasing concerns about their security and robustness (He et al. 2016; Liu et al. 2021). It is now well-established that DNNs are highly vulnerable to adversarial attacks (Szegedy et al. 2014a), wherein imperceptible, carefully crafted perturbations are added to clean inputs to generate adversarial examples that can mislead the model into producing incorrect outputs (Huang and Shen 2025).
@@
To address this issue, adversarial training (AT) (Madry et al. 2019) has been proposed, which retrains classifiers using adversarial examples. However, AT suffers from high computational cost and poor generalization to unseen threats, limiting its applicability in real-world adversarial defense scenarios.
@@
In contrast, adversarial purification (AP) has emerged as a compelling alternative due to its plug-and-play nature—requiring no retraining of the classifier—and its stronger generalization capabilities. AP methods utilize generative models as a preprocessing step to transform adversarial examples into purified ones, which are then fed into the classifier for accurate predictions. The recent advances in diffusion models (Ho, Jain, and Abbeel 2020) have further propelled the development of AP. These models learn to transform simple distributions into complex data distributions through a forward noising and reverse denoising process. Crucially, this iterative denoising mechanism aligns well with the goal of removing adversarial perturbations, making diffusion models a natural fit for AP tasks (Nie et al. 2022).
@@
However, existing diffusion-based purification approaches suffer from a critical limitation: they require multiple iterative denoising steps, resulting in prohibitively slow inference, which severely restricts their use in latency-sensitive applications such as autonomous driving (Chi et al. 2024) and industrial manufacturing (Wang et al. 2025). Moreover, most of these methods rely on a key assumption that the distributions of clean and adversarial samples converge after a certain number of forward diffusion steps. This allows the use of pretrained diffusion models, originally designed for generative tasks, to purify adversarial samples. However, this assumption only holds when the diffusion time horizon is sufficiently large. Empirical evidence from DiffPure suggests that excessive diffusion steps can lead to significant loss of semantic content, rendering accurate reconstruction of clean images infeasible.
@@
In this paper, we propose Diffusion Bridge Distillation for Purification (DBLP), a novel framework designed to simultaneously address the two key limitations of existing diffusion-based adversarial purification methods: low inference efficiency and detail degradation. At its core, DBLP introduces a noise-bridged alignment strategy within the Latent Consistency Model (Luo et al. 2023a), effectively bridging adversarial noise and clean targets during the consistency distillation process to better align with the purification objective. By leveraging noise bridge distillation, DBLP enables direct recovery of clean samples from diffused adversarial inputs using an ODE solver. To further mitigate detail loss caused by fewer denoising steps, we introduce adaptive semantic enhancement, a lightweight yet effective conditioning mechanism that utilizes multi-scale pyramid edge maps to capture fine-grained structural features. These semantic priors are injected into inference to enhance content preservation. DBLP achieves SOTA robust accuracy across multiple benchmark datasets while substantially reducing inference latency, requiring only 0.1 seconds per sample, thus making real-time adversarial purification feasible without compromising visual quality.
@@
  • We propose DBLP, a novel diffusion-based adversarial purification framework that significantly accelerates inference while preserving high visual quality and improving purification performance.
@@
Related Work
Adversarial Training
@@
Adversarial training is a prominent defense strategy against adversarial attacks (Goodfellow, Shlens, and Szegedy 2015), which enhances model robustness by retraining the model on perturbed adversarial examples (Lau et al. 2023). A substantial body of research has demonstrated its efficacy in adversarial defense. Notable methods include min-max optimization framework (Madry et al. 2018), TRADES which balances robustness and accuracy via a regularized loss (Zhang et al. 2019), and techniques like local linearization (Qin et al. 2019) and mutual information optimization (Zhou et al. 2022). Despite its strong robustness, adversarial training suffers from several notable drawbacks. It often generalizes poorly to unseen attacks (Laidlaw, Singla, and Feizi 2021), and it incurs significant computational overhead due to the necessity of retraining the entire model. Moreover, it typically leads to a degradation in clean accuracy (Wong, Rice, and Kolter 2020).
Adversarial Purification
@@
Adversarial purification represents an alternative and effective defense strategy against adversarial attacks that circumvents the need for retraining the model. The core idea is to employ generative models to pre-process adversarially perturbed images, yielding purified versions that are subsequently fed into the classifier. Early efforts in this domain leveraged GANs (Samangouei, Kabkab, and Chellappa 2018) or score-based matching techniques (Yoon, Hwang, and Lee 2021; Song et al. 2021) to successfully restore adversarial images. DiffPure (Nie et al. 2022) advanced this with diffusion models, inspiring follow-ups like adversarially guided denoising (Wang et al. 2022; Wu, Ye, and Gu 2022), improved evaluation frameworks (Lee and Kim 2023), gradient-based purification (Zhang et al. 2023a), dual-phase guidance (Song et al. 2024), and adversarial diffusion bridges (Li et al. 2025). Despite their promising results, these methods exhibit certain limitations. Many approaches rely on auxiliary classifiers, which often compromise generalization performance. Others involve iterative inference procedures that are computationally intensive and time-consuming, thereby limiting their practicality in real-time or resource-constrained scenarios.
Diffusion Models
@@
Diffusion models (Ho, Jain, and Abbeel 2020), originally introduced to enhance image generation capabilities, have since demonstrated remarkable success across various domains, including video synthesis (Ho et al. 2022) and 3D content generation (Luo and Hu 2021). As a class of score-based generative models, diffusion models operate by progressively corrupting images with Gaussian noise in the forward process, and subsequently generating samples by denoising in the reverse process (Huang and Tang 2025). Given a pre-defined forward trajectory $\{\mathbf{x}_{t}\}_{t\in[0,T]}$, indexed by a continuous time variable $t$, the forward process can be effectively modeled using a widely adopted stochastic differential equation (SDE) (Karras et al. 2022):
$\mathrm{d}\mathbf{x}_{t}=\boldsymbol{\mu}(\mathbf{x}_{t},t)\mathrm{d}t+\sigma(t)\mathrm{d}\mathbf{w}_{t},$ (1)
@@
where $\boldsymbol{\mu}(\mathbf{x}_{t},t)$ and $\sigma(t)$ denote the drift and diffusion coefficients, respectively, while $\{\mathbf{w}_{t}\}_{t\in[0,T]}$ represents a standard $d$-dimensional Brownian motion. Let $p_{t}(\mathbf{x})$ denote the marginal distribution of $\mathbf{x}_{t}$ at time $t$, and $p_{\mathrm{data}}(\mathbf{x})$ represent the distribution of the original data, then $p_{0}(\mathbf{x})=p_{\mathrm{data}}(\mathbf{x})$.
@@
Remarkably, (Song et al. 2021) established the existence of an ordinary differential equation (ODE), referred to as the Probability Flow (PF) ODE, whose solution trajectories share the same marginal probability densities $p_{t}(\mathbf{x})$ as those of the forward SDE:
$\mathrm{d}\mathbf{x}_{t}=\left[\boldsymbol{\mu}\left(\mathbf{x}_{t},t\right)-\frac{1}{2}\sigma(t)^{2}\nabla\log p_{t}(\mathbf{x}_{t})\right]\mathrm{d}t.$ (2)
@@
For sampling, a score model $s_{\boldsymbol{\phi}}(\mathbf{x},t)\approx\nabla\log p_{t}(\mathbf{x})$ is first trained via score matching to approximate the gradient of the log-density at each time step. This learned score function is then substituted into Equation (2) to obtain an empirical estimate of the PF ODE:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\boldsymbol{\mu}\left(\mathbf{x}_{t},t\right)\mathbf{x}_{t}-\frac{1}{2}\sigma(t)^{2}s_{\boldsymbol{\phi}}(\mathbf{x}_{t},t).$ (3)
Preliminaries
Figure 1: The overview structure of our DBLP. An adversary perturbs a clean image $\mathbf{x}$ with noise $\boldsymbol{\epsilon}_{a}$ into an adversarial example $\mathbf{x}^{a}$, we first encode it into the latent space using the encoder $\mathcal{E}$ to obtain the latent representation $\mathbf{z}$, followed by noise injection as defined in Equation (11). During training, we adopt a modified LCM-LoRA framework to perform noise bridge consistency distillation on the diffusion model, and employ a leapfrog ODE solver to accelerate sampling. During inference, we introduce adaptive semantic enhancement, using the weighted fusion of pyramid edge maps as a semantic-preserving condition to guide the purification process. The final purified image $\mathbf{x}_{\mathrm{pur}}$ is then recovered via the decoder $\mathcal{D}$.
Problem Formulation
@@
Adversarial attacks were first introduced by (Szegedy et al. 2014b), who revealed the inherent vulnerability of neural networks to carefully crafted perturbations. An adversarial example $\mathbf{x}_{\mathrm{adv}}$ is visually and numerically close to a clean input $\mathbf{x}$, yet it is deliberately designed to mislead a classifier $\mathrm{C}$ into assigning it to an incorrect label, rather than the true class $y_{\mathrm{true}}$, formally expressed as:
@@
Consistency Models
@@
The long inference time of diffusion models is a well-known limitation, prompting the introduction of the Consistency Model (Song et al. 2023), which enables the sampling process to be reduced to just a few steps, or even a single step. It proposes learning a direct mapping from any point $\mathbf{x}_{t}$ along the PF ODE trajectory $\{\mathbf{x}_{t}\}_{t\in[0,T]}$ back to its starting point, referred to as the consistency function, denoted as $\boldsymbol{f}:(\mathbf{x}_{t},t)\mapsto\mathbf{x}_{\epsilon}$, where $\mathbf{x}_{\epsilon}$ represents the starting state at a predefined small positive value $\epsilon$. The self-consistency property of this function can be formalized as:
$\boldsymbol{f}(\mathbf{x}_{t},t)=\boldsymbol{f}(\mathbf{x}_{t^{\prime}},t^{\prime})\quad\forall t,t^{\prime}\in[0,T].$ (6)
@@
The goal of the consistency model $\boldsymbol{f_{\theta}}$ is to estimate the underlying consistency function $\boldsymbol{f}$ by enforcing the self-consistency property. The model $\boldsymbol{f_{\theta}}$ can be parameterized as:
$\boldsymbol{f_{\theta}}(\mathbf{x},t)=c_{\mathrm{skip}}(t)\mathbf{x}+c_{\mathrm{out}}(t)F_{\boldsymbol{\theta}}(\mathbf{x},t),$ (7)
@@
where $c_{\mathrm{skip}}(t)$ and $c_{\mathrm{out}}(t)$ are differentiable functions. To satisfy the boundary condition $\boldsymbol{f}(\mathbf{x}_{\epsilon},\epsilon)=\mathbf{x}_{\epsilon}$, we have $c_{\mathrm{skip}}(\epsilon)=1$ and $c_{\mathrm{out}}(\epsilon)=0$.
@@
Building on this, the Latent Consistency Model (Luo et al. 2023a) extends the consistency model to the latent space using an auto-encoder (Rombach et al. 2022). In this setting, the consistency function conditioned on $\mathbf{c}$ is defined as $\boldsymbol{f_{\theta}}:(\mathbf{z}_{t},\mathbf{c},t)\mapsto\mathbf{z}_{\epsilon}$. To fully leverage the capabilities of a pretrained text-to-image model, LCM parameterizes the consistency model as:
@@
$\boldsymbol{f_{\theta}}(\mathbf{z},\mathbf{c},t)=c_{\mathrm{skip}}(t)\mathbf{z}+c_{\mathrm{out}}(t)\left(\frac{\mathbf{z}-\sigma_{t}\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{z},\mathbf{c},t)}{\alpha_{t}}\right),$ (8)
@@
LCM-LoRA (Luo et al. 2023b) proposes distilling LCM using LoRA, significantly reducing the number of trainable parameters and thereby greatly decreasing training time and computational cost.
Methodology
Overall Framework
@@
In this work, we aim to accelerate the purification backbone using a consistency distillation-inspired approach. Noting that the starting and ending points of the ODE trajectory respectively contain and exclude adversarial perturbations, we propose Noise Bridge Distillation in Section Noise Bridge Distillation to explicitly align the purification objective.
@@
To achieve acceleration, we leverage the Latent Consistency Model with LoRA-based distillation and introduce a leapfrog ODE solver for efficient sampling. During inference, as detailed in Section Adaptive Semantic Enhanced Purification, we propose Adaptive Semantic Enhancement, which fuses pyramid edge maps into a semantic-preserving condition to guide the diffusion model toward effective purification.
Noise Bridge Distillation

### v2

@@ 1 Introduction
Deep neural networks (DNNs) have achieved remarkable success across a wide range of tasks in recent years. However, their widespread deployment has raised increasing concerns about their security and robustness He et al. (2016); Liu et al. (2021). It is now well-established that DNNs are highly vulnerable to adversarial attacks Szegedy et al. (2014a), wherein imperceptible, carefully crafted perturbations are added to clean inputs to generate adversarial examples that can mislead the model into producing incorrect outputs Huang & Shen (2025).
@@
To address this issue, adversarial training (AT) Madry et al. (2019) has been proposed, which retrains classifiers using adversarial examples. However, AT suffers from high computational cost and poor generalization to unseen threats, limiting its applicability in real-world adversarial defense scenarios.
@@
In contrast, adversarial purification (AP) has emerged as a compelling alternative due to its stronger generalization capabilities, and its plug-and-play nature, requiring no classifier retraining. AP methods utilize generative models as a preprocessing step to transform adversarial examples into purified ones, which are then fed into the classifier. The recent advances in diffusion models Ho et al. (2020) have further propelled the development of AP. These models learn to transform simple distributions into complex data distributions through a forward noising and reverse denoising process. Crucially, this iterative denoising mechanism aligns well with the goal of removing adversarial perturbations, making diffusion models a natural fit for AP tasks Nie et al. (2022).
@@
However, existing diffusion-based purification approaches suffer from a critical limitation: they require multiple iterative denoising steps, resulting in prohibitively slow inference, which severely restricts their use in latency-sensitive applications such as autonomous driving Chi et al. (2024) and industrial manufacturing Wang et al. (2025). Moreover, most of these methods rely on a key assumption that the distributions of clean and adversarial samples converge after a certain number of forward diffusion steps. This allows the use of pretrained diffusion models, originally designed for generative tasks, to purify adversarial samples. However, this assumption only holds when the diffusion time horizon is sufficiently large. Empirical evidence from DiffPure Nie et al. (2022) suggests that excessive diffusion steps can lead to significant loss of semantic content, rendering accurate reconstruction of clean images infeasible.
@@
In this paper, we propose Diffusion Bridge Distillation for Purification (DBLP), a novel framework designed to simultaneously address the two key limitations of existing diffusion-based adversarial purification methods: low inference efficiency and detail degradation. At its core, DBLP introduces a noise-bridged alignment strategy within the Latent Consistency Model Luo et al. (2023a), effectively bridging adversarial noise and clean targets during the consistency distillation process to better align with the purification objective. By leveraging noise bridge distillation, DBLP enables direct recovery of clean samples from diffused adversarial inputs using an ODE solver. To further mitigate detail loss caused by fewer denoising steps, we introduce adaptive semantic enhancement, a lightweight yet effective conditioning mechanism that utilizes multi-scale pyramid edge maps to capture fine-grained structural features. These semantic priors are injected into inference to enhance content preservation. DBLP achieves SOTA robust accuracy across multiple benchmark datasets while substantially reducing inference latency, requiring only 0.2 seconds per sample, thus making real-time adversarial purification feasible without compromising visual quality.
@@
  • We propose DBLP, a novel diffusion-based adversarial purification framework that significantly accelerates inference while improving purification performance and visual quality.
@@
2 Related Work
2.1 Adversarial Training
@@
Adversarial training is a prominent defense strategy against adversarial attacks Goodfellow et al. (2015), which enhances model robustness by retraining the model on perturbed adversarial examples Lau et al. (2023). A substantial body of research has demonstrated its efficacy in adversarial defense. Notable methods include min-max optimization framework Madry et al. (2018), TRADES which balances robustness and accuracy via a regularized loss Zhang et al. (2019), and techniques like local linearization Qin et al. (2019) and mutual information optimization Zhou et al. (2022). Despite its strong robustness, adversarial training suffers from several notable drawbacks. It often generalizes poorly to unseen attacks Laidlaw et al. (2021), and it incurs significant computational overhead due to the necessity of retraining the entire model. Moreover, it typically leads to a degradation in clean accuracy Wong et al. (2020).
2.2 Adversarial Purification
@@
Adversarial purification represents an alternative and effective defense strategy against adversarial attacks that circumvents the need for retraining the model. The core idea is to employ generative models to pre-process adversarially perturbed images, yielding purified versions that are subsequently fed into the classifier. Early efforts in this domain leveraged GANs Samangouei et al. (2018) or score-based matching techniques Yoon et al. (2021); Song et al. (2021) to successfully restore adversarial images. DiffPure Nie et al. (2022) advanced this with diffusion models, inspiring follow-ups like adversarially guided denoising Wang et al. (2022); Wu et al. (2022), improved evaluation frameworks Lee & Kim (2023), gradient-based purification Zhang et al. (2023a), dual-phase guidance Song et al. (2024), and adversarial diffusion bridges Li et al. (2025). Despite their promising results, these methods exhibit certain limitations. Many approaches rely on auxiliary classifiers, which often compromise generalization performance. Others involve iterative inference procedures that are computationally intensive and time-consuming, thereby limiting their practicality in real-time or resource-constrained scenarios.
2.3 Diffusion Models
@@
Diffusion models Ho et al. (2020), originally introduced to enhance image generation capabilities, have since demonstrated remarkable success across various domains, including video synthesis Ho et al. (2022) and 3D content generation Luo & Hu (2021). As a class of score-based generative models, diffusion models operate by progressively corrupting images with Gaussian noise in the forward process, and subsequently generating samples by denoising in the reverse process Huang & Tang (2025). Given a pre-defined forward trajectory $\{\mathbf{x}_{t}\}_{t\in[0,T]}$, indexed by a continuous time variable $t$, the forward process can be effectively modeled using a widely adopted stochastic differential equation (SDE) Karras et al. (2022):
$\mathrm{d}\mathbf{x}_{t}=\bm{\mu}(\mathbf{x}_{t},t)\mathrm{d}t+\sigma(t)\mathrm{d}\mathbf{w}_{t},$ (1)
@@
where $\bm{\mu}(\mathbf{x}_{t},t)$ and $\sigma(t)$ denote the drift and diffusion coefficients, respectively, while $\{\mathbf{w}_{t}\}_{t\in[0,T]}$ represents a standard $d$-dimensional Brownian motion. Let $p_{t}(\mathbf{x})$ denote the marginal distribution of $\mathbf{x}_{t}$ at time $t$, and $p_{\mathrm{data}}(\mathbf{x})$ represent the distribution of the original data, then $p_{0}(\mathbf{x})=p_{\mathrm{data}}(\mathbf{x})$.
@@
Remarkably, Song et al. (2021) established the existence of an ordinary differential equation (ODE), referred to as the Probability Flow (PF) ODE, whose solution trajectories share the same marginal probability densities $p_{t}(\mathbf{x})$ as those of the forward SDE:
$\mathrm{d}\mathbf{x}_{t}=\left[\bm{\mu}\left(\mathbf{x}_{t},t\right)-\frac{1}{2}\sigma(t)^{2}\nabla\log p_{t}(\mathbf{x}_{t})\right]\mathrm{d}t.$ (2)
@@
For sampling, a score model $s_{\bm{\phi}}(\mathbf{x},t)\approx\nabla\log p_{t}(\mathbf{x})$ is first trained via score matching to approximate the gradient of the log-density at each time step. This learned score function is then substituted into Equation equation 2 to obtain an empirical estimate of the PF ODE:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\bm{\mu}\left(\mathbf{x}_{t},t\right)-\frac{1}{2}\sigma(t)^{2}s_{\bm{\phi}}(\mathbf{x}_{t},t).$ (3)
3 Preliminaries
3.1 Problem Formulation
@@
Adversarial attacks were first introduced by Szegedy et al. (2014b), who revealed the inherent vulnerability of neural networks to carefully crafted perturbations. An adversarial example $\mathbf{x}_{\mathrm{adv}}$ is visually and numerically close to a clean input $\mathbf{x}$, yet it is deliberately designed to mislead a classifier $\mathrm{C}$ into assigning it to an incorrect label, rather than the true class $y_{\mathrm{true}}$, formally expressed as:
@@
3.2 Consistency Models
Figure 1: The overview structure of our DBLP. An adversary perturbs a clean image $\mathbf{x}$ with noise $\bm{\epsilon}_{a}$ into an adversarial example $\mathbf{x}^{a}$, we first encode it into the latent space using the encoder $\mathcal{E}$ to obtain the latent representation $\mathbf{z}$, followed by noise injection as defined in Equation equation 11. During training, we adopt a modified LCM-LoRA framework to perform noise bridge consistency distillation on the diffusion model, and employ a leapfrog ODE solver to accelerate sampling. During inference, we introduce adaptive semantic enhancement, using the weighted fusion of pyramid edge maps as a semantic-preserving condition to guide the purification process. The final purified image $\mathbf{x}_{\mathrm{pur}}$ is then recovered via the decoder $\mathcal{D}$.
@@ 3.2 Consistency Models
The long inference time of diffusion models is a well-known limitation, prompting the introduction of the Consistency Model Song et al. (2023), which enables the sampling process to be reduced to just a few steps, or even a single step. It proposes learning a direct mapping from any point $\mathbf{x}_{t}$ along the PF ODE trajectory $\{\mathbf{x}_{t}\}_{t\in[0,T]}$ back to its starting point, referred to as the consistency function, denoted as $\bm{f}:(\mathbf{x}_{t},t)\mapsto\mathbf{x}_{\epsilon}$, where $\mathbf{x}_{\epsilon}$ represents the starting state at a predefined small positive value $\epsilon$. The self-consistency property of this function can be formalized as:
$\bm{f}(\mathbf{x}_{t},t)=\bm{f}(\mathbf{x}_{t^{\prime}},t^{\prime})\quad\forall t,t^{\prime}\in[0,T].$ (6)
@@
The goal of the consistency model $\bm{f_{\theta}}$ is to estimate the underlying consistency function $\bm{f}$ by enforcing the self-consistency property. The model $\bm{f_{\theta}}$ can be parameterized as:
$\bm{f_{\theta}}(\mathbf{x},t)=c_{\mathrm{skip}}(t)\mathbf{x}+c_{\mathrm{out}}(t)F_{\bm{\theta}}(\mathbf{x},t),$ (7)
@@
where $c_{\mathrm{skip}}(t)$ and $c_{\mathrm{out}}(t)$ are differentiable functions. To satisfy the boundary condition $\bm{f}(\mathbf{x}_{\epsilon},\epsilon)=\mathbf{x}_{\epsilon}$, we have $c_{\mathrm{skip}}(\epsilon)=1$ and $c_{\mathrm{out}}(\epsilon)=0$.
@@
Building on this, the Latent Consistency Model Luo et al. (2023a) extends the consistency model to the latent space using an auto-encoder Rombach et al. (2022). In this setting, the consistency function conditioned on $\mathbf{c}$ is defined as $\bm{f_{\theta}}:(\mathbf{z}_{t},\mathbf{c},t)\mapsto\mathbf{z}_{\epsilon}$. To fully leverage the capabilities of a pretrained text-to-image model, LCM parameterizes the consistency model as:
@@
$\bm{f_{\theta}}(\mathbf{z},\mathbf{c},t)=c_{\mathrm{skip}}(t)\mathbf{z}+c_{\mathrm{out}}(t)\left(\frac{\mathbf{z}-\sigma_{t}\hat{\bm{\epsilon}}_{\theta}(\mathbf{z},\mathbf{c},t)}{\alpha_{t}}\right),$ (8)
@@
LCM-LoRA Luo et al. (2023b) proposes distilling LCM using LoRA, significantly reducing the number of trainable parameters and thereby greatly decreasing training time and computational cost.
4 Methodology
4.1 Overall Framework
@@
In this work, we aim to accelerate the purification backbone using a consistency distillation-inspired approach. Noting that the starting and ending points of the ODE trajectory respectively contain and exclude adversarial perturbations, we propose Noise Bridge Distillation in Section 4.2 to explicitly align the purification objective.
@@
To achieve acceleration, we leverage the Latent Consistency Model with LoRA-based distillation and introduce a leapfrog ODE solver for efficient sampling. During inference, as detailed in Section 4.3, we propose Adaptive Semantic Enhancement, which fuses pyramid edge maps into a semantic-preserving condition to guide the diffusion model toward effective purification.
4.2 Noise Bridge Distillation